# Index Terms:

### v1

@@ V Proposed Channel Estimation Algorithm
where $\bm{\Lambda}$ is the Lagrangian dual variable and $\rho>0$ is a hyperparameter. In each iteration, the ADMM algorithm performs the following operations: {mdframed} ADMM for Solving (10):
${\bm{z}}^{\ell}=\underset{{\bm{z}}}{\arg\min}\left\|{\cal G}({\bm{z}})-\left({\bm{H}}^{\ell-1}-\frac{\bm{\Lambda}^{\ell-1}}{\rho}\right)\right\|_{\rm{F}}^{2},$ (29a)
${\bm{H}}^{\ell}=\underset{{\bm{H}}}{\arg\min}-\log p({\bm{r}}|{\bm{h}})+\frac{\rho}{2}\left\|{\bm{H}}-{\cal G}\left({\bm{z}}^{\ell}\right)-\frac{\bm{\Lambda}^{\ell-1}}{\rho}\right\|_{\rm{F}}^{2},$ (29b)
$\bm{\Lambda}^{\ell}=\bm{\Lambda}^{\ell-1}+\rho\left({\cal G}({\bm{z}}^{\ell})-{\bm{H}}^{\ell}\right),$ (29c)
where $\ell$ denotes the iteration number. As mentioned, the MLE problem (10) is challenging because of the manifold structure (represented by ${\cal G}(\bm{z})$). Under the ADMM update rule, estimating $\bm{z}$ from ${\cal G}(\bm{z})$ is now only present in the subproblem (29a), having a relatively “clean” least squares-based denoising form. Next, we describe how to solve the subproblems (29a) and (29b), respectively.
@@ V-A Modified RELAX for Subproblem (29a)
The subproblem (29a) can be viewed as an HR problem, if one regards (${\bm{H}}^{\ell-1}-\bm{\Lambda}^{\ell-1}/\rho$) as a noisy multi-dimensional harmonic signal [27, 30]. This is because, similar to HR, Eq. (29a) also aims to recover the frequency components $(\bm{\theta},\bm{\beta})$ and amplitudes $\bm{\beta}$ from the noisy two-dimensional signal $({\bm{H}}^{\ell-1}-\bm{\Lambda}^{\ell-1}/\rho)$. Noisy multi-dimensional HR poses a challenging nonconvex optimization problem. Nonetheless, as a classical task in array processing, many sophisticated HR algorithms exist. We propose to employ the celebrated RELAX method [24] to tackle (29a). Note that the original RELAX algorithm essentially discretizes the spatial domain (i.e., $[-\pi,\pi]$) and estimates $\theta_{i}$ and $\phi_{i}$ over discrete spatial grids. This is not desired in CSI recovery. Hence, we also propose a simple modification step to the RELAX iterations as remedy.
@@
where we have $\beta_{i}=\nicefrac{{({\bm{a}}_{t}(\phi_{i})\otimes{\bm{a}}_{r}(\theta_{i}))^{\rm{H}}\bm{\zeta}_{i}}}{{MN}}$ and the matricization function “${\rm mat}(\cdot)$” is the inverse operation of “${\rm vec}(\cdot)$”. The original RELAX algorithm approximates the solution of (33) using 2D-FFT. RELAX is widely recognized as an effective method for handling 2D HR [27]. Nonetheless, the 2D-FFT step works over a discretized array manifold, which means that the accuracy of RELAX is limited by the discretization resolution (i.e., the size of the DFT matrix used). To pull the RELAX-found solutions back to the continuous space, we further refine the solution by running gradient ascent for solving (33); that is,
@@
where $f_{i,\eta}^{\rm HR}(\phi_{i},\theta_{i})=|{\bm{a}}_{r}(\theta_{i})^{\rm{H}}\operatorname{mat}(\bm{\zeta}_{i}){\bm{a}}_{t}(\phi_{i}))|^{2}$. The gradient updates in (34) follows the 2D-FFT step for solving (33). The modified RELAX (Mod-RELAX) algorithm is summarized in Appendix A.
@@ V-B EM Algorithm for Subproblem (29b)
Many algorithms can handle the subproblem in (29b), e.g., the proximal gradient algorithm proposed in the context of quantized compressive sensing [47]. We propose to employ an expectation-maximization (EM) approach that showed empirical success in handling similar objective functions as in (29b) [48, 49]. The EM algorithm consists of the following two steps:
@@
To summarize, the $\bm{z}$-update and $\bm{H}$-update in our ADMM algorithm in (29) are carried out by Mod-RELAX and EM, respectively. The algorithm in (29) was for solving Problem (10), i.e., the MLE under Scheme 1. To apply the ADMM algorithm to solving Problem (15) under Scheme 2, one can simply replace ${\bm{r}}$, ${\bm{A}}$, ${\bm{h}}$ and $\sigma_{v}$ by ${\bm{r}}$, ${\bm{D}}$, $\check{{\bm{h}}}$ and $\check{\sigma}$ in (15), respectively; also note that the correspondence between ${\bm{h}}$ and ${\bm{H}}$ needs to be replaced by the correspondence between $\check{{\bm{h}}}$ and ${\bm{H}}$. We refer to our algorithm in (29) as RELAX and EM-assisted alternating direction method of multipliers (REDEEM). Fig. 3 shows the sharp contrast between the convergence characteristics of the REDEEM algorithm and directly applying gradient descent (GD) to Problem (10).
@@ VI Numerical Results
VI-A Simulation Settings
@@ VI-A1 Data Generation
In the simulations, the BS and the user are assumed to be equipped with ULAs, with the inter-antenna spacing $d=\lambda/2$. The channel is generated according to the double directional channel model in (4). The AoD and AoA angles $\{\phi_{i}\}_{i},\{\theta_{i}\}_{i}$ are randomly sampled from the uniform distribution on $(-\frac{\pi}{2},\frac{\pi}{2})$. Each complex path loss $\beta_{i}$ is generated with its amplitude $|\beta_{i}|$ and angle $\angle\beta_{i}$ uniformly sampled from $(0.5,1)$ and $(0,2\pi)$, respectively. The quantizer has uniform quantization intervals. Our results are averaged over 1,000 trials with randomly generated data in each trial.
@@ VI-A2 Performance Metric
Following the literature [14, 50, 51, 52], we use the averaged normalized mean squared error (NMSE) to measure the CSI reconstruction performance of all algorithms. Given the ground-truth channel matrix ${\bm{H}}^{\natural}$ and the estimate $\widehat{{\bm{H}}}$, the NMSE is defined as follows:
$\text{NMSE}=\mathbb{E}\left[\frac{\|{\bm{H}}^{\natural}-\widehat{{\bm{H}}}\|_{\rm{F}}^{2}}{\|{\bm{H}}^{\natural}\|_{\rm{F}}^{2}}\right].$ (38)
@@ VI-A3 Baselines
We compare our proposed methods with a diverse set of baselines. First, we consider approaches that use limited feedback and relegate most computational burden of the channel estimation to the BS side. This includes the sparse regression based MLE (SparseMLE) [14], and EMMP [51]. Note that the original SparseMLE method uses its own compression strategy (where the compression matrix is essentially a Kronecker product of $\bm{S}^{T}$ and a Gaussian compression matrix). We also implement a version of SparseMLE that uses the estimation-and-then-compression strategy of Scheme 1. We refer to this variant as SparseMLE (Scheme 1). Second, as a reference, we also present the results of methods that carry out heavy computations at the BS for estimating the key parameters from ${\bm{H}}$. This includes OMP-SQ [14] and UE-HR that performs HR on $\bm{H}$ at the UE side; in our experiments, UE-HR employs RELAX [24] for HR. These results are used to provide insights into the gap between the accuracy of full-precision measurement-based estimations and those relying on quantized measurements.
@@ VI-A4 Algorithm Settings
VI-B Results
(a) $N=32$, $M=16$, $K=6$, 2-bit quantization.
(b) $N=32$, $M=24$, $K=8$, 3-bit quantization.
Figure 4: Recovery performance under different $Q$’s.
Fig. 4 shows the channel estimation performance under various numbers of feedback bits. Fig. 4(a) and (b) presents the results under $Q=2$ bits and $Q=3$ bits, respectively. It is seen that both of the proposed schemes achieve promising channel estimation accuracy. In particular, Fig. 4 (a) shows that when 500 bits are used, Schemes 1 and 2 reach NMSEs of around $2\times 10^{-2}$ and $4\times 10^{-2}$, respectively, while the baselines all admit NMSEs larger than $10^{-1}$. Note that if one feeds back the whole $\bm{H}$ with full precision (i.e., using 32 bits to represent a real value), 32,768 bits would be needed. Using 500 bits costs 1.5% of the overhead.
@@ VI-B Results
Figure 5: Recovery performance under different feedback dimensions ($R$ and $T$ in Schemes 1 and 2, respectively); $N=32$, $M=16$, $K=6$. Solid and dashed lines represent the 3-bit quantization case and the 2-bit quantization case, respectively.
Fig. 6 and Fig. 7 compare the proposed schemes with baselines that estimate the channel parameters at the UE, namely, OMP-SQ and UE-HR. The former presents the NMSE performance and the latter shows the “preparation time” performance, which is the processing time needed at UE before sending the feedback bits. The preparation time reflects the complexity of UE operations, e.g., matched filtering in (5), Gaussian compression in (7), dithered quantization in (8), or the rank-1 projection of our Scheme 2 in (11). For methods that estimate $(\bm{\theta},\bm{\beta},\bm{\phi})$ at the UE, the preparation time also includes the time for executing the channel estimation algorithms. In general, one does not wish to spend long preparation time at the UE side due to the fast changing nature of wireless channels. It is seen that both OMP-SQ and UE-HR yield favorable NMSE performance. Nonetheless, they attain low NMSEs at the cost of sharply increased computational burden and the preparation time at the UE side. On the other hand, our proposed method only requires light process at the UE side—that is, when $Q\times R=600$ our Schemes 1 and 2 only cost 0.0419% and 0.00419% of UE-HR’s preparation time to reach a reasonable NMSE of $1\times 10^{-2}$, respectively.
Figure 6: Performance comparison between the proposed and UE-based methods; $N=32$, $M=16$, $K=6$, $Q=3$. Fig. 7 shows the feedback time required for each of these methods.
Figure 7: Feedback preparation time at UE; $M=16$, $K=6$, 3-bit quantization, 900 feedback bits.
Fig. 8 shows the CSI recovery performance under different SNRs, where we fix the compression dimension $R=300$ and use $Q=3$. It is seen that all the considered algorithms get better as SNR increases. Our proposed schemes show a significant performance margin than existing designs—the NMSEs are at least 40% lower than those of the baselines when SNR5absent5$\geq$ 5≥ 5dB.
Fig. 9 shows the impact of increasing the number of $N$, i.e., the number of transmitter antennas. The NMSEs of the proposed schemes are almost unaffected under this test setting. In contrast, the performance of other methods degrades faster as the antenna size increases.
Figure 8: Sensitivity to SNR of received pilot signals; $N=32$, $M=16$, $K=6$, $Q=3$, 900 feedback bits.
Figure 9: Recovery performance as the number of antennas grows; $M=16$, $K=6$, $Q=3$, 900 feedback bits.

### v2

@@ V Proposed Channel Estimation Algorithm
where $\bm{\Lambda}$ is the Lagrangian dual variable and $\rho>0$ is a hyperparameter. In each iteration, the ADMM algorithm performs the following operations: ADMM for Solving (10): ${\bm{z}}^{\ell}=\underset{{\bm{z}}}{\arg\min}\left\|{\cal G}({\bm{z}})-\left({\bm{H}}^{\ell-1}-\frac{\bm{\Lambda}^{\ell-1}}{\rho}\right)\right\|_{\rm{F}}^{2},$ (29a) ${\bm{H}}^{\ell}=\underset{{\bm{H}}}{\arg\min}-\log p({\bm{r}}|{\bm{h}})+\frac{\rho}{2}\left\|{\bm{H}}-{\cal G}\left({\bm{z}}^{\ell}\right)-\frac{\bm{\Lambda}^{\ell-1}}{\rho}\right\|_{\rm{F}}^{2},$ (29b) $\bm{\Lambda}^{\ell}=\bm{\Lambda}^{\ell-1}+\rho\left({\cal G}({\bm{z}}^{\ell})-{\bm{H}}^{\ell}\right),$ (29c) where $\ell$ denotes the iteration number. As mentioned, the MLE problem (10) is challenging because of the manifold structure (represented by ${\cal G}(\bm{z})$). Under the ADMM update rule, estimating $\bm{z}$ from ${\cal G}(\bm{z})$ is now only present in the subproblem (29a), having a relatively “clean” least squares-based denoising form. Next, we describe how to solve the subproblems (29a) and (29b), respectively.
@@ V-A Modified RELAX for Subproblem (29a)
The subproblem (29a) can be viewed as an HR problem, if one regards (${\bm{H}}^{\ell-1}-\bm{\Lambda}^{\ell-1}/\rho$) as a noisy multi-dimensional harmonic signal [27, 30]. This is because, similar to HR, Eq. (29a) also aims to recover the frequency components $(\bm{\theta},\bm{\phi})$ and amplitudes $\bm{\beta}$ from the noisy two-dimensional signal $({\bm{H}}^{\ell-1}-\bm{\Lambda}^{\ell-1}/\rho)$. Noisy multi-dimensional HR poses a challenging nonconvex optimization problem. Nonetheless, as a classical task in array processing, many sophisticated HR algorithms exist. We propose to employ the celebrated RELAX method [24] to tackle (29a). Note that the original RELAX algorithm essentially discretizes the spatial domain (i.e., $\pi,\pi]$) and estimates $\theta_{i}$ and $\phi_{i}$ over discrete spatial grids. This is not desired in CSI recovery. Hence, we also propose a simple modification step to the RELAX iterations as remedy.
@@
where we have $\beta_{i}=\nicefrac{{({\bm{a}}_{t}^{*}(\phi_{i})\otimes{\bm{a}}_{r}(\theta_{i}))^{\rm{H}}\bm{\zeta}_{i}}}{{MN}}$ and the matricization function ${\rm mat}(\cdot)$” is the inverse operation of ${\rm vec}(\cdot)$”. The original RELAX algorithm approximates the solution of (33) using 2D-FFT. RELAX is widely recognized as an effective method for handling 2D HR [27]. Nonetheless, the 2D-FFT step works over a discretized array manifold, which means that the accuracy of RELAX is limited by the discretization resolution (i.e., the size of the DFT matrix used). To pull the RELAX-found solutions back to the continuous space, we further refine the solution by running gradient ascent for solving (33); that is,
@@
where $f_{i,\eta}^{\rm HR}(\phi_{i},\theta_{i})=|{\bm{a}}_{r}(\theta_{i})^{\rm{H}}\operatorname{mat}(\bm{\zeta}_{i}){\bm{a}}_{t}(\phi_{i})|^{2}$. The gradient updates in (34) follows the 2D-FFT step for solving (33). The modified RELAX (Mod-RELAX) algorithm is summarized in Appendix A.
@@ V-B EM Algorithm for Subproblem (29b)
Many algorithms can handle the subproblem in (29b), e.g., the proximal gradient algorithm proposed in the context of quantized compressive sensing [47]. We propose to employ an expectation-maximization (EM) approach that showed empirical success in handling similar objective functions as in (29b) [48, 49]. The EM algorithm is based on optimization of the following variational upper bound:
$\begin{split}&\leavevmode\nobreak\ -\log\left[\Phi\left(\frac{\bar{b}[r_{i}]-{\bm{a}}_{i}^{\top}{\bm{h}}^{j}}{\sigma_{v}}\right)-\Phi\left(\frac{\underline{b}[r_{i}]-{\bm{a}}_{i}^{\top}{\bm{h}}^{j}}{\sigma_{v}}\right)\right]\\ \leq&\leavevmode\nobreak\ \frac{1}{2\sigma_{v}^{2}}|g_{i}-{\bm{a}}_{i}^{T}{\bm{h}}|^{2}+c,\end{split}$
where
$g_{i}=\leavevmode\nobreak\ {\bm{a}}_{i}^{\top}{\bm{h}}^{j}-\sigma\frac{\phi\left(\frac{\bar{b}[r_{i}]-{\bm{a}}_{i}^{\top}\bar{{\bm{h}}}^{j}}{\sigma_{v}}\right)-\phi\left(\frac{\underline{b}[r_{i}]-{\bm{a}}_{i}^{\top}\bar{{\bm{h}}}^{j}}{\sigma_{v}}\right)}{\Phi\left(\frac{\bar{b}[r_{i}]-{\bm{a}}_{i}^{\top}\bar{{\bm{h}}}^{j}}{\sigma_{v}}\right)-\Phi\left(\frac{\underline{b}[r_{i}]-{\bm{a}}_{i}^{\top}\bar{{\bm{h}}}^{j}}{\sigma_{v}}\right)}$ (35)
for any $\bar{{\bm{h}}}$, and $c$ is a constant independent of $\bm{h}$}. The equality holds when $\bar{{\bm{h}}}={\bm{h}}$.
The EM algorithm consists of the following two steps: updating the upperbound followed by maximization with respect to this upperbound.
@@
To summarize, the $\bm{z}$-update and $\bm{H}$-update in our ADMM algorithm in (29) are carried out by Mod-RELAX and EM, respectively. The algorithm in (29) was for solving Problem (10), i.e., the MLE under Scheme 1. To apply the ADMM algorithm to solving Problem (15) under Scheme 2, one can simply replace $\bm{r}$}, $\bm{A}$}, $\bm{h}$} and $\sigma_{v}$ by $\bm{r}$}, $\bm{D}$}, $\check{{\bm{h}}}$ and $\check{\sigma}$ in (15), respectively; also note that the correspondence between $\bm{h}$} and $\bm{H}$} needs to be replaced by the correspondence between $\check{{\bm{h}}}$ and $\bm{H}$}. We refer to our algorithm in (29) as RELAX and EM-assisted alternating direction method of multipliers (REDEEM). Fig. 3 shows the sharp contrast between the convergence characteristics of the REDEEM algorithm and directly applying gradient descent (GD) to Problem (10). We also present detailed per-iteration complexity analysis for REDEEM in Appendix D.
@@ VI Numerical Results
In this section, we will first present basic simulations using the signal model (2). Then, we will test the methods under more challenging scenarios generated by the DeepMIMO toolbox [50].
VI-A Basic Simulations
@@ VI-A1 Data Generation
In the simulations, the BS and the user are assumed to be equipped with ULAs, with the inter-antenna spacing $d=\lambda/2$. The channel is generated according to the double directional channel model in (4). The AoD and AoA angles $\{\phi_{i}\}_{i},\{\theta_{i}\}_{i}$ are randomly sampled from the uniform distribution on $(-\frac{\pi}{2},\frac{\pi}{2})$. Each complex path loss $\beta_{i}$ is generated with its amplitude $|\beta_{i}|$ and angle $\angle\beta_{i}$ uniformly sampled from $(0.5,1)$ and $(0,2\pi)$, respectively. The quantizer has uniform quantization intervals. The min and max values of 1000 data samples, treated as historic CSI samples, are used as the min and max quantization boundaries in our experiments. The dithering noise level $\sigma_{v}$ (and $\check{\sigma}_{v}$) is set to $25\%$ of the maximal signal amplitude. Our results are averaged over 1,000 trials with randomly generated data in each trial. Unless otherwise noted, all the experiments in Sec. VI-A use above settings.
@@ VI-A2 Performance Metric
Following the literature [14, 51, 52, 53], we use the averaged normalized mean squared error (NMSE), and average beamforming gain (BGain) to measure the CSI reconstruction performance of all algorithms. Given the ground-truth channel matrix $\bm{H}$$^{\natural}$ and the estimate $\widehat{{\bm{H}}}$, the NMSE and BGain is defined as follows:
$\text{NMSE}=\mathbb{E}\left[\frac{\|{\bm{H}}^{\natural}-\widehat{{\bm{H}}}\|_{\rm{F}}^{2}}{\|{\bm{H}}^{\natural}\|_{\rm{F}}^{2}}\right],$ (39)
$\text{BGain}=\mathbb{E}\left[{\rm tr}(\widehat{{\bm{H}}}^{H}{\bm{H}})^{2}/\|\widehat{{\bm{H}}}\|_{F}^{2}\right].$ (40)
@@ VI-A3 Baselines
We compare our proposed methods with a diverse set of baselines. First, we consider approaches that use limited feedback and relegate most computational burden of the channel estimation to the BS side. This includes the sparse regression based MLE (SparseMLE) [14], and EMMP [52]. Note that the original SparseMLE method uses its own compression strategy (where the compression matrix is essentially a Kronecker product of $\bm{S}^{\top}$ and a Gaussian compression matrix). We also implement a version of SparseMLE that uses the estimation-and-then-compression strategy of Scheme 1. We refer to this variant as SparseMLE (Scheme 1). We apply the same dithering technique and noise level for all the benchmarking methods. Second, as a reference, we also present the results of methods that carry out heavy computations at the UE for estimating the key parameters from $\bm{H}$}. This includes OMP-SQ [14] and UE-HR that performs HR on $\bm{H}$ at the UE side; in our experiments, UE-HR employs RELAX [24] for HR. These results are used to provide insights into the gap between the accuracy of full-precision measurement-based estimations and those relying on quantized measurements.
@@ VI-A4 Algorithm Settings
VI-A5 Results
(a) $(M,K,Q)=(16,6,2)$.
(b) $(M,K,Q)=(24,8,3)$.
Figure 4: NMSE Performance under different $Q$’s when $N=32$.
(a) $(M,K,Q)=(16,6,2)$.
(b) $(M,K,Q)=(24,8,3)$.
Figure 5: BGain performance under different $Q$’s when $N=32$.
Fig. 4 and Fig. 5 show the NMSE and BGain performance, respectively, under various numbers of feedback bits. Fig. 4 (a) and (b) present the NMSE results under $Q=2$ bits and $Q=3$ bits, respectively. It is seen that both of the proposed schemes achieve promising channel estimation accuracy. In particular, Fig. 4 (a) shows that when 500 bits are used, Schemes 1 and 2 reach NMSEs of around $2\times 10^{-2}$ and $4\times 10^{-2}$, respectively, while the baselines all admit NMSEs larger than $10^{-1}$. Fig. 5 (b) shows that when $Q=3$ and 500 bits are used for feedback, our methods are able to attain a BGain value larger than $430$, while other methods are below around $360$. Note that if one feeds back the whole $\bm{H}$ with full precision (i.e., using 32 bits to represent a real value), 32,768 bits would be needed. Using 500 bits costs 1.5% of the overhead.
Fig.  6 shows the NMSE performance under a $900$-bit feedback budget as $Q$ changes. Note that both of our schemes enjoy growing accuracy when $Q$ increases from 2 to 4 and attain their best performance at $4$-bit quantization. This is because increased $Q$ makes brings more fine-grained quantization interval. However, when $Q>5$, allocating more bits to the quantizer leaves too few measurements ($R$ for Scheme $1$ and $T$ for Scheme $2$) to for feedback, hurting the overall performance. This presents an interesting tradeoff for the system designers to balance with.
Figure 6: NMSE performance under different $Q$’s with a fixed 900 bits feedback budget; $N=32$, $M=16$ $K=6$.
@@ VI-A5 Results
Figure 7: NMSE performance under different number of measurements ($R$ and $T$ for Scheme 1 and 2); $N=32$, $M=16$, $K=6$; Solid line: $Q=3$; Dash line: $Q=2$.
Fig. 8 (a) and Fig. 8 (b) compare the proposed schemes with baselines that estimate the channel parameters at the UE, namely, OMP-SQ and UE-HR. The former presents the NMSE performance and the latter shows the “preparation time (PT)” performance, which is the processing time needed at UE before sending the feedback bits. The preparation time reflects the complexity of UE operations, e.g., Gaussian compression in (7), dithered quantization in (8), or the rank-1 projection of our Scheme 2 in (11). For methods that estimate $(\bm{\theta},\bm{\beta},\bm{\phi})$ at the UE, the preparation time also includes the time for executing the channel estimation algorithms (e.g., least squares). In general, one does not wish to spend long preparation time at the UE side due to the fast changing nature of wireless channels. It is seen that both OMP-SQ and UE-HR yield favorable NMSE performance. Nonetheless, they attain low NMSEs at the cost of sharply increased computational burden and the preparation time at the UE side. On the other hand, our proposed method only requires light process at the UE side—that is, when $Q\times R=600$ our Schemes 1 and 2 only cost 0.0419% and 0.00419% of UE-HR’s preparation time to reach a reasonable NMSE of $1\times 10^{-2}$, respectively. In particular, to achieve the same channel estimation accuracy as Scheme 1, Scheme 2 reduces the computational complexity at the UE side without significantly increasing the feedback overhead. This may help reduce the overall delay at the BS in estimating the downlink channel.
(a) NMSE of the methods.
(b) PT at UE; 900 feedback bits.
Figure 8: Performance of the proposed methods and the UE-based methods; $N=32$, $M=16$, $K=6$, $Q=3$.
[…]
Fig. 10 validates the our recoverability theorems. In particular, Fig. 10 (a) shows the impact of increasing the number of $N$, i.e., the number of transmitter antennas. The NMSEs of the proposed schemes are almost unaffected under this test setting. In contrast, the performance of other methods degrades faster as the antenna size increases. The insensitivity of our methods to the growing $N$ is consistent with our theorems: The performance of our method is not affected by $M$ or $N$ when $R$ is larger than $\Omega(K)$. Fig. 10 (b) shows that the algorithm-obtained MSE values indeed decays with a rate roughly at the predicted order, namely, $O(1/\sqrt{R})$ and $O(1/\sqrt{T})$ for Schemes 1 and 2, respectively.
(a) NMSE vs SNR; $N=32$, $M=16$, $K=6$, $Q=3$, 900 feedback bits.
(b) NMSE vs $R$ and $T$ when $|\beta_{i}|$’s are sampled from [0.01,1].
Figure 9: Robustness evaluation with respect to SNR of received signals and high dynamic range of path loss
(a) NMSE v.s. NN; $M=16$, $K=6$, $Q=3$, 900 feedback bits.
(b) MSE vs $R$ and $T$.
Figure 10: Theorem validation: insensitivity to $N$ and MSE decaying rate when $R$ and $T$ grow.
@@
VI-B DeepMIMO-based Simulations
VI-B1 Data Generation
We generate the channel matrix by using the DeepMIMO-v2 tool [50, 54] that uses ray-tracing to simulate the real-world channels. We present our results in three different scenarios provided by DeepMIMO: Outdoor 1 Blockage Scenario (O1), Boston5G Scenario (B5), and Indoor 3 Scenario (I3).
For each scenario, DeepMIMO uses several predefined BS locations, identified using labels such as “BS1” and “BS2”. For the O1 scenario, we use BS3, and for both B5 and I3, we select BS1. The exact positions of these BS are shown in their corresponding scenario figures; see [50] and the illustration in https://www.deepmimo.net/scenarios/. The UEs are placed into the environment space uniformly at random. For both the BS and UE, we use the same antenna configuration and interelement spacing as in Section VI-A, along with identical quantization intervals and dithering noise levels. All experiments here employ $N=32$ transmit antennas, $M=16$ receive antennas, $K=6$ propagation paths, and $3$-bit quantization.
Compared to our basic simulations, DeepMIMO creates more complex and challenging scenarios. DeepMIMO labels each channel with its Line-of-Sight (LOS) status, denoting whether an unobstructed direct path exists between the BS and UE exists. We refer to those with a clear direct path as With-Line-of-Sight (WLOS) channels, and those without any direct path as Non-Line-of-Sight (NLOS) channels. The DeepMIMO uses the ray-tracing model to generates channel whose combined path loss statistics are listed in Table. I; “combined” here means an overall characterization of all path losses. It is seen that the path losses generated by DeepMIMO have a wide dynamic range, presenting challenging channel estimation scenarios.
Table I: Combined path loss characteristics for the DeepMIMO scenarios; provided [50].
Scenario Max (dB) Min (dB) Mean $\pm$ Std (dB)
O1 $151.41$ $58.93$ $86.51\pm 8.88$
B5 $133.03$ $60.61$ $93.02\pm 13.77$
I3 $79.25$ $45.18$ $53.89\pm 3.21$
VI-B2 Results
Fig. 12 compares the NMSE performance of the methods for the O1 scenario. Fig. 12 (a) shows the performance for WLOS paths, whereas 12 (b) shows for NLOS paths. It is seen that our methods have promising channel estimation results compared to the baselines—consistent with the results observed in the basic simulations. Similarly, Figs. 13 and 14 showthat our proposed methods work well under channel configurations under the B5 and I3 scenarios, respectively.
(a) WLOS
(b) NLOS
Figure 12: Performance under various numbers of measurements under the O1 scenario; $N=32$, $M=16$, $K=6$, $Q=3$.
(a) WLOS
(b) NLOS
Figure 13: Performance under various numbers of measurements under the B5 scenario; $N=32$, $M=16$, $K=6$, $Q=3$.
(a) WLOS
(b) NLOS
Figure 14: Recovery performance under different feedback dimension for I3 scenario; $N=32$, $M=16$, $K=6$, 33-bit quantization.